OpenStudy (anonymous):

18% of 90 is what number

OpenStudy (akashdeepdeb):

What this means is that! In 100 it is 18 parts! Then in 90 it is how many parts? So we find 18/100 of 90 Like this..... 18 * 90/100 = 16.2 Got it? :)

OpenStudy (anonymous):

\[\frac{ 18 }{ 100 } \times 90 = \frac{ 18 \times 90 }{ 100 } = \frac{ 18 \times 9 }{ 10 }\]

OpenStudy (anonymous):

\[ = \frac{ 172 }{ 10 } = 17.2\]